# Early-Exit Graph Neural Networks

### v1

@@ Appendix G Extended Related Work
Early-Exit Neural Networks (General). Early-exit methods in standard deep learning reduce inference cost by attaching auxiliary classifiers at intermediate layers. Training paradigms include deeply supervised nets [Lee et al., 2014], layer-wise pretraining [Hettinger et al., 2017], and independent exit training [Venkataramani et al., 2015]. At inference, early termination is governed by confidence- or entropy-based thresholds [Wang et al., 2017, Xin et al., 2020, Baccarelli et al., 2020, Xin et al., 2021] or threshold-agnostic strategies [Pomponi et al., 2022, Ju et al., 2021]. These techniques inspire our threshold-free, jointly trainable exit mechanism tailored to graph-structured inputs and obviate the need for hand-tuned confidence thresholds. Despite this, we do not draw inspiration from any specific method, but we recognize the benefit that early-exit can have in graph learning tasks.

### v2

@@ Appendix G Extended Related Work
Early-Exit in Neural Networks (General). Early-exit has been widely explored in standard deep learning to reduce inference cost by attaching auxiliary classifiers to intermediate layers. Training paradigms include deeply supervised nets [Lee et al., 2014], layer-wise pretraining [Hettinger et al., 2017], and independent exit training [Venkataramani et al., 2015]. Inference termination typically depends on confidence thresholds [Wang et al., 2017, Xin et al., 2020, Baccarelli et al., 2020, Xin et al., 2021] or threshold-free approaches [Pomponi et al., 2022, Ju et al., 2021]. These works inspire our threshold-free, jointly trainable exit mechanism, tailored to graph-structured data and obviating the need for hand-tuned thresholds.